Problem 3-15 Journal Entries; T-Accounts; Financial Statements (LO3-1, LO3-2, LO3-3, LO3-4] Froya Fabrikker A/S of Bergen, NoComplete this question by entering your answers in the tabs below. Req 1 Reg 2 Reg 3 Req 4A Req 4B Reg 5 Prepare journal entrReg 1 Reg 2 Req3 Reg 4A Reg 4B Reg 5 Post your entries to T-accounts. (Dont forget to enter the beginning inventory balancesComplete this question by entering your answers in the tabs below. Req 1 Reg 2 Req 3 Req 4A Req 4B Req 5 Prepare a schedule oJournal entry worksheet < 1 Record the entry to close any balance in the manufacturing overhead account to cost of goods soldComplete this question by entering your answers in the tabs below. Req 1 Reg 2 Req3 Req 4A Req 4B Req 5 Prepare a schedule ofComplete this question by entering your answers in the tabs below. Reg 1 Reg 2 Reg 3 Req 4A Req 4B Req 5 Prepare an income st
Answer #1

Answer with explanation:

1.

Transaction General Journal Debit $ 290,000 Credit $ a. Raw Materials Inventory Accounts Payable 290,000 b. 275,000 Work in P

2.

Raw Materials Inventory $48,000 b) 290,000 Beg. bal. a) 275,000 Beg. bal. b) d) i) Work-in-process Inventory $39,0001). 950,0

3.

Froya Fabrikker A/S Schedule of Cost of Goods Manufactured Direct materials: Beginning raw materials inventory $48,000 Add: P

4.A)

Credit Debit $8,300 Transaction General Journal 1 Manufacturing overhead Cost of goods sold (To record overapplied overhead)

Predetermined overhead rate = ($395,600/ 920 hours) × 970 direct labor hours = $417,100


Actual overhead = 69,300+108,000+72,000+67,500+92,000 = $408,800

Overapplied or underapplied overhead = $417,100 - $408,800 = $8,300 overapplied.
( When the actual overhead is more than applied, it is called overapplied)